If 800g of a radioactive substance are present initially and 8 years later only 450g remain, how much of the substance will be present after 16 years?
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



Over the space of 8 years, the amount remaining decreases by a factor of 450/800 = 9/16.

After another 8 years, the amount will again decrease by the same factor.

ANSWER: (9/16)*450 = 253 1/8

Since radioactive decay is a statistical process, it should be okay to round the answer to 253g.


Answer by ikleyn(52794)

.

On radioactive decay, see the lesson
    - Radioactive decay problems
in this site.

Find there numerous solved similar and different problems and learn the methods of their solution.

Do not miss this opportunity  ( ! )